OpenStudy (anonymous):

Which statement describes the error made in the following solution process? k minus seven equals negative eight The addition calculation was incorrect. Subtraction should have been used instead of addition. Seven was added to the left side of the equation instead of eight. Eight was added to the right side of the equation instead of seven.

OpenStudy (anonymous):

I think we're missing the actual solution process here.

OpenStudy (anonymous):

k-7=-8 +7 +8 k=0

OpenStudy (anonymous):

When you solve an equation, you have to do the same operation to both sides, to ensure that the two sides remain equal. In this process, 7 was added to the left side, but 8 was added to the right side. This is the error, and is the fourth option.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome, glad I could help